A mercury drop of radius 1 mm is sprayed into 10^3


drops of equal sizes. Compute


the energy expended. Surface tension of mercury = 3.5 × 10^−2 N/m.


1
Expert's answer
2021-03-08T08:17:41-0500
"\Delta E=4\pi r^2(n^{\frac{1}{3}}-1)T\\\Delta E=4\pi (0.001)^2(1000^{\frac{1}{3}}-1)0.035=4\cdot10^{-6}\ J"